INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM FLUORIDE IN THE NEAR IN-FRARED

\underline{LEAH C. O’BRIEN}, RACHEL A. HARRIS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

A new electronic transition of PtF has been recorded using intracavity laser absorption spectroscopy. This is the first reported spectroscopic observation of PtF. The results of the analysis will be presented. The gas phase PtF molecules were produced using a platinum-lined hollow cathode in an argon-based electric discharge with a small amount of SF_{6}.